### v2

# Homophily Within and Across Groups

Abstract
Homophily—the tendency to connect with similar others—plays a central role in shaping network structure and function. It is often treated as a uniform, global parameter, independent from other structural features. Here, we propose a maximum-entropy framework that decomposes homophily into contributions within and across groups, with the stochastic block model emerging as a special case. Our exponential-family formulation, parameterized by group size, fits real-world social networks well and allows homophily to be captured with a single parameter per group size. This decomposition shows that aggregate metrics often obscure group-level variation. We also find that the scale-dependent distribution of homophily has a significant impact on network percolation, influencing epidemic thresholds, the spread of ideas or behaviors, and the effectiveness of intervention strategies. Ignoring this heterogeneity can lead to distorted conclusions about connectivity and dynamics in complex systems.
Epidemic Spreading, Homophily
I Introduction
Homophily—the tendency to interact with similar others—shapes both the structure and function of networks [1, 2, 3, 4, 5]. In social systems, people tend to connect with others who share attributes such as sex, age, race, occupation, sexual preferences, socioeconomic background, or vaccination status [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. Assortative mixing occurs in various settings, ranging from friendships and workplace ties to online platforms, peer influence networks [17, 18, 19, 20], and even graph neural networks [21, 22, 23, 24]. It shapes interactions and governs processes like contagion, diffusion, and coordination [25, 26, 27, 28, 29]. In healthcare settings, the close contact between healthcare workers and vulnerable patients can alter outbreak patterns, as homophily interacts with variations in infectiousness and susceptibility, reshaping epidemic dynamics [30, 31, 32, 33, 34, 35, 36, 37].
Homophily varies by interaction type and tie strength [38, 39, 40]. Strong ties tend to form within close-knit groups, while weak ties more often bridge distant or dissimilar clusters [41, 39]. Networks of citations and trade partnerships typically exhibit dual mixing patterns, combining dense local collaboration with broad cross-group exchange [42, 43]. Gender homophily, in particular, is sensitive to both medium and group size. On Instagram, women show strong same-gender preferences in comments, while men are largely neutral [44]. In face-to-face proximity networks, women tend to form homophilic triads, whereas men favor same-sex dyads [45]. In contrast, the teen-focused Spanish social network Tuenti shows the opposite: women sustain stronger same-gender dyads, and men more often form all-male triads than expected under null models [46].
These observations indicate that homophily is influenced by both group composition and size [47], highlighting its variability across social scales [48]. Small, cohesive groups often display high local homophily, whereas long-range connections mix more randomly, producing distinct within- and across-group connectivity essential for understanding collective behavior [49]. Nevertheless, many network models simplify homophily into a single, uniform parameter—essentially a summary statistic that captures only the network’s average mixing pattern [50, 25, 14, 51, 52, 53, 54]. Epidemiological studies emphasize the importance of this complexity, demonstrating that high within-group homophily can elevate epidemic thresholds, while increased cross-group interactions typically lower them [6, 55, 56, 57, 58, 59]. Thus, to improve our understanding of network formation and dynamics, we must model homophily as a heterogeneous quantity across social scales.
We introduce a network model in which groups, modeled as cliques, exhibit homophily that varies with group size. Using a maximum-entropy framework, the model captures homophilic structure beyond pairwise links and distinguishes between within- and across-group connections. This reveals scale-dependent patterns often overlooked by standard measures and aligns well with empirical studies. Our results show that such multi-scale homophily can raise or lower the percolation threshold, depending on how intra- and inter-group links are distributed. These findings offer new tools for understanding epidemic thresholds, designing interventions, and modeling the spread of information or behavior in social systems.
The remainder of the paper is structured as follows. Section II describes our network model and demonstrates how it encodes multi-scale homophily. Section III validates the model using real-world data, illustrating how within- and across-group mixing patterns emerge under the same modeling assumptions. Section IV examines how multi-scale homophily influences percolation thresholds and network connectivity, with implications for epidemic control and other contagion processes.
Figure 1: Constructing a Homophilic Clique Network with Eight Red and Seven Blue Nodes. With two colors, each $c$-clique can appear in $c+1$ different compositions, distributed according to ${F}_{c}$. (a-c) The network $G$ is constructed by merging multiple clique layers—specifically, 2-, 3-, and 4-cliques—where each layer $G_{c}$ contains cliques of size $c$ but shares the same set of nodes. While the nodes remain fixed, each layer differs in clique size $c$ and composition distribution ${F}_{c}$. The accompanying histograms show the frequency of each $c$-clique type in the network $G_{c}$. In each layer, $M_{c}$ groups of $c$ nodes are sampled with replacement from the available colored nodes and converted into $c$-cliques. For example, $G_{3}$ consists of seven sampled 3-cliques, while $G_{2}$ is constructed similarly, but with its own parameters $M_{2}$ and ${F}_{2}$. (d) All layers are then merged to form the final network: $G=\bigcup G_{c}$. (e) The maximum-entropy clique composition distribution ${F}_{5}$, derived from Eq. (8), is illustrated for cliques of size $c=5$ at homophily levels $h_{5}=0.2$ and $h_{5}=0.6$, across several values of the red-node fraction $n_{\mathrm{r}}$. For clarity, we omit the corresponding networks generated from these distributions. More details in Sec. II.
II Model Description
We introduce a network model that captures homophily at different social scales. The model generates an undirected, sparse graph $G$, formed as the union of multiple subnetworks $\{G_{c}\}$, all sharing the same set of nodes. Each subnetwork $G_{c}$ consists of cliques of size $c$ (groups). Every node in the network has a discrete attribute—for example, a categorical label indicating membership. We begin with a binary labeling of nodes as red or blue, extending to multiple classes in Sec. V.4. As shown in Fig.1, the network consists of $N$ nodes, with $N_{\mathrm{r}}$ red and $N_{\mathrm{b}}=N-N_{\mathrm{r}}$ blue. Their corresponding proportions are $n_{\mathrm{r}}=N_{\mathrm{r}}/N$ and $n_{\mathrm{b}}=1-n_{\mathrm{r}}$. All subnetworks $G_{c}$ share the same node set; that is, red and blue labels are consistent across layers.
We classify the cliques within each subnetwork $G_{c}$ based on the number of red or blue nodes they contain. A $c$-clique can be classified into one of $c+1$ types, that are specified by the number $i$ of red nodes included. We denote by $F_{c,i}$ the proportion of $c$-cliques in $G_{c}$ that contain exactly $i$ red nodes and $c-i$ blue nodes. To ensure that the overall fraction of red and blue nodes is preserved in the large-network limit, the distributions $\{F_{c,i}\}$ must satisfy
$\sum_{i=0}^{c}\frac{i}{c}\,F_{c,i}=n_{\mathrm{r}},\qquad\sum_{i=0}^{c}\frac{c-i}{c}\,F_{c,i}=n_{\mathrm{b}}.$ (1)
To construct each subnetwork $G_{c}$, we independently sample $M_{c}$ cliques of size $c$ from the total population of nodes according to the distribution $\{F_{c,i}\}$. Fig. 1(a-c) illustrates the construction of three subnetworks with different clique-type distributions. In this framework, $G_{1}$ is a trivial graph of isolated colored nodes, and $G_{2}$ is a two-block stochastic block model [60]. The final network $G$, as shown in Fig. 1d, is obtained by merging all subnetworks;
$G=\bigcup_{c}G_{c}.$ (2)
Since cliques are sampled independently, setting $M_{c}=\mathcal{O}(N)$ ensures that each layer $G_{c}$ remains sparse and that the fraction of overlapping links vanishes in the large-network limit. Given a prescribed average degree $\langle d\rangle$, controlling $M_{c}$ allows us to regulate the number of links from each clique size to the overall network. In practice, we sample fewer large cliques, reflecting empirical observations: larger cliques tend to be rarer in real-world networks due to cognitive and social constraints—a phenomenon known as “schisming” [61, 62]. A more detailed discussion regarding finiteness and empirical networks is provided in Sec. V.2.
II.1 Group Homophily Values
A standard global measure for quantifying how nodes with categorical attributes (such as color) tend to connect is the Coleman index, also known as the assortativity coefficient [63, 64, 65]. It is defined as
$h=\frac{\sum_{k}(e_{kk}-n_{k}^{2})}{1-\sum_{k}n_{k}^{2}}=\frac{e_{\mathrm{rr}}+e_{\mathrm{bb}}-n_{\mathrm{r}}^{2}-n_{\mathrm{b}}^{2}}{1-n_{\mathrm{r}}^{2}-n_{\mathrm{b}}^{2}},$ (3)
where $e_{\mathrm{rr}}$ and $e_{\mathrm{bb}}$ are the fractions of red–red and blue–blue links, and $n_{k}$ is the fraction of nodes of color $k$. The numerator in Eq. 3 corresponds to modularity, while $h$ is modularity normalized by its maximum value [66, 67]. By construction, $|h|\leq 1$: values near 1 signal strong homophily or segregation; $h\approx 0$ indicates random mixing; and negative values imply heterophily [68].
In our framework, cliques are sampled independently, and each node participates in a Poisson-distributed number of cliques. Under these assumptions, the Coleman index (3) can be decomposed into contributions from each clique layer $G_{c}$, giving
$h=\sum_{c\geq 2}\alpha_{c}\,h_{c},$ (4)
where $h_{c}$ denotes the homophily of $G_{c}$, and $\alpha_{c}$ is the fraction of links contributed by layer $G_{c}$. Consider a network $G$ constructed from two layers, $G_{2}$ and $G_{6}$. Here, $h_{2}$ quantifies across-group homophily and $h_{6}$ captures within-group homophily, assuming the groups are defined by the 6-cliques. The total homophily becomes $h=\alpha_{2}h_{2}+\alpha_{6}h_{6},$ with each layer’s influence determined by its corresponding $\alpha_{c}$. For instance, $\alpha_{2}=\alpha_{6}=0.5$ gives equal weight to both layers, whereas $\alpha_{2}>\alpha_{6}$ results in across-group homophily dominating, and vice versa. In general, the weights $\alpha_{c}$ can be computed as
$\alpha_{c}=\frac{c(c-1)}{N\langle d\rangle}M_{c}.$ (5)
Each $h_{c}$ can further be decomposed into the weighted average of homophily contributions $h_{c,i}$ from all $c$-cliques with $i$ red nodes:
$h_{c}=\sum_{i=0}^{c}h_{c,i}\,F_{c,i}.$ (6)
$h_{c,i}$ measures the deviation of such a $c$-clique’s composition from random mixing and is defined as
$h_{c,i}=\frac{\tfrac{\binom{i}{2}+\binom{c-i}{2}}{\binom{c}{2}}-n_{\mathrm{r}}^{2}-n_{\mathrm{b}}^{2}}{1-n_{\mathrm{r}}^{2}-n_{\mathrm{b}}^{2}}.$ (7)
This equation defines homophily at the clique level, following the Coleman approach. It is worth noting that Eq. 3 can also be decomposed into contributions at the node or neighborhood level, where “node congruity” measures each node’s share of the overall homophily [69, 70, 48]. However, the formalism presented above provides sufficient measures for this study.
II.2 Maximum Entropy Distribution
Since many different $F_{c,i}$ distributions can produce the same overall homophily $h_{c}$, we use a maximum entropy approach to identify the least-biased distribution consistent with the desired group homophily $h_{c}$ and the red and blue node fractions $n_{\mathrm{r}}$ and $n_{\mathrm{b}}$. Specifically, we choose $F_{c,i}$ to maximize entropy subject to two constraints: (i) the average fraction of red nodes in cliques equals $n_{\mathrm{r}}$ (Eq. (1)); and (ii) the resulting clique-layer homophily is $h_{c}$ (Eq. (6)). Since the network contains only two colors, satisfying constraint (i) automatically ensures that the blue-node fraction $n_{\mathrm{b}}=1-n_{\mathrm{r}}$ is preserved, so only one constraint is needed to fix both. The resulting distribution takes the exponential-family form:
$F_{c,i}=\frac{1}{Z}\exp{(\theta_{\mathrm{r}}i+\theta_{h_{c}}h_{c,i})},$ (8)
where $Z$ is the partition function normalizing over all $c$-clique compositions. The parameters $\theta_{\mathrm{r}}$ and $\theta_{h_{c}}$ emerge naturally as Lagrange multipliers in our maximum-entropy formulation [71]: $\theta_{\mathrm{r}}$ sets the fraction of red nodes, while $\theta_{h_{c}}$ tunes the homophily level $h_{c}$. These parameters are commonly referred to as conjugate couplings in statistical mechanics. This approach naturally accommodates networks with different degrees of mixing and handles both symmetric and skewed attribute distributions. Fig. 1e shows three examples from this versatile family of distributions.
Analogous to spin models on hyperedges [72, 73, 74], our framework can be viewed as a canonical ensemble, where clique homophily and color composition act as soft constraints—similar to inverse-temperature-like parameters in a spin system defined on a hypergraph. Each clique configuration corresponds to a microstate, weighted by a Boltzmann factor $F\sim e^{-\mathcal{H}}$. In this formulation, the Hamiltonian $\mathcal{H}$ depends on the full color composition of each clique. A $c$-clique thus represents a higher-order interaction among the spins (i.e., node attributes) of its $c$ members.
This two-color framework easily generalizes to multiple categorical attributes and even multidimensional homophily, allowing more attributes to be considered simultaneously. Sections V.4 and V.5 detail these generalizations explicitly.
III Results
To assess how well our model captures homophilic structure in real-world networks, we analyze six datasets with categorical node attributes: the Slovak social networking site Pokec [75], the Copenhagen Networks Study (CPH) [76], the music-sharing platform Last.fm [77], a mobile phone call network [78, 79], the face-to-face proximity dataset SocioPatterns [80, 45], and 100 Facebook friendship networks [81]. These datasets, summarized in Table 1, span a broad spectrum of social interaction types—from digital connections to physical encounters.
For each empirical network $G$, we extract all maximal cliques and group them by size to form distinct layers $\{G_{c}\}$. As shown in Fig. 1(a–d), this decomposition yields a family of subgraphs, each capturing interactions at a specific clique size. Some links appear in multiple layers, but this redundancy has a negligible effect on measured homophily. We fit the maximum-entropy model (8) to the observed clique-type distributions by maximizing the likelihood, thereby estimating the parameters $\{\theta\}$. For details on edge duplication and its impact, see Sec. V.2.
Our maximum-entropy approach closely matches the empirical distributions of clique compositions, as illustrated in Fig. 2a. Similar to Fig. 1, it compares observed histograms of $F_{c}$ across several empirical networks with the model fits from Eq. (8). With only two parameters—the expected homophily $h_{c}$ and the fraction of nodes of one type—the model accurately captures the full distribution of clique compositions $F_{c}$ in our empirical data. This means that $h_{c}$ values contain all the information about the complete distributions $F_{c}$ of our data, and it is enough to report the $h_{c}$ values to fully describe how the homophily is distributed within and across groups of different sizes.
Although these networks differ in global properties, such as degree distributions that do not necessarily align with those generated by our model, the method still effectively reproduces the patterns of homophily. This robustness stems from the fact that the model is directly constrained by empirical clique compositions or homophily values, rather than by assumptions about the degree sequence or other global network characteristics.
Total Coleman indices range from mildly homophilic ($h=0.1$ in the Last.fm network and $h=0.2$ in the University of Oklahoma Facebook network), to randomly mixed ($h=0.0$ in the CPH and SocioPatterns networks), and mildly heterophilic ($h=-0.2$ on the Pokec platform). However, these aggregate measures conceal substantial variation at the clique level. We observe diverging trends across datasets. In the mildly homophilic networks, homophily increases with clique size: Last.fm exhibits random mixing for dyads ($h=0$) but stronger homophily in larger cliques ($h=0.3$); Oklahoma shows heterophily among bridging links ($h=-0.2$) and homophily only in larger groups. In contrast, CPH and SocioPatterns—despite appearing randomly mixed overall—display the opposite pattern: bridging links are homophilic, but larger cliques are heterophilic.
The Pokec network offers a particularly striking example of a pattern behind an average index. Its Coleman index $h=-0.2$ suggests mild heterophily, yet dyads ($G_{2}$) exhibit strong cross-sex preference with $h_{2}=-0.6$. As clique size increases from 2 to 10, this intermingling turns into homophily. This transformation is not a minor refinement—it fundamentally changes the interpretation of the network: pairwise links reflect Pokec’s dating-oriented function, while larger cliques reflect its broader role as a social media platform where, at social scales, birds of a feather flock together [1].
Dataset Link Type $n_{\mathrm{r}}$ (%) $h$ $N$
Mobile Phone Phone Calls 55.1 $0.1$ 2,173,030
Pokec Friendship 46.6 $-0.2$ 383,943
Last.fm Friendship 68.3 $0.1$ 188,672
Oklahoma FB Friendship 49.3 $0.2$ 16,245
CPH Proximity, Calls, etc. 78.1 $0.0$ 779
SocioPatterns Proximity 57.4 $0.0$ 115
Table 1: Summary of datasets used in this study. The variable $n_{\mathrm{r}}$ represents the relative size of the red (male) group. The Coleman index $h$ is for overall sex homophily.
Figure 2: Empirical and Theoretical Distributions of Clique Types and Homophily in Real-world Networks mentioned in Sec. III. (a) Each column compares empirical clique-type distributions (histograms) with the corresponding maximum-entropy distributions (dashed lines, computed using Eq. (8)), showing strong agreement and validating the model’s accuracy in capturing observed clique compositions across real-world datasets. Red represents males and blue, females. (b) Variation in homophily values $h_{c}$ across clique sizes (up to $c=10$) for several empirical networks, highlighting distinct group interaction patterns based on sex attributes. Error bars represent 95% confidence intervals obtained via bootstrap resampling. (c–e) Average homophily trends across clique sizes for Facebook friendship networks from 100 U.S. institutions, with attributes grouped by: (c) Sex (two categories), (d) Student status (two categories), and (e) Class year (12 categories). Gray background curves represent individual institutions, while colored lines indicate the average trends. Lighter colors reflect a greater number of contributing institutions, emphasizing both inter-institutional variability and how homophily depends on the attribute under consideration.
Facebook data from 100 U.S. universities and colleges offers a socially comparable set of environments with rich attribute annotations. Altenburger et al. [82] previously showed that users tend to form highly sex-homogeneous friendship circles—often all-male or all-female—even though these preferences largely cancel out in aggregate, producing weak overall homophily. Fig. 2(c–e) illustrates how homophily varies with clique size when nodes are grouped by (c) sex, (d) student status, and (e) class year. Each attribute captures different social dynamics, leading to distinct homophily profiles. In Fig. 2c, sex-based homophily increases monotonically with clique size, suggesting that sex-based clustering intensifies in larger groups. In contrast, Fig. 2d shows an average trend with a medium level of homophily at the link level and only a slight dependency on group size.
Our framework can analyze more complex scenarios involving multiple non-binary attributes. Fig. 2e reveals a clearly non-monotonic pattern for class year in the Facebook network: homophily rises in smaller cliques but levels off as groups grow, likely reflecting increased mixing between cohorts in broader social circles. Notably, similar multiscale mixing patterns have been reported by Peel et al.[48], who demonstrated that even when the network exhibits modest overall homophily, certain subgroups (such as first-year students) display markedly higher local homophily.
These observations shed a new light on findings by Traud et al. [81], who showed that communities in university Facebook networks tend to cluster by class year. Our analysis reveals that smaller cliques strongly reflect graduation-year homophily, but this pattern weakens in larger groups, where other attributes—such as sex or student status—become more prominent. Similar scale-dependent patterns are well documented in the broader sociological literature. Homophily is shaped by multiple, context-dependent mechanisms [1]; academic performance influences how students reorganize their friendships [83]; and ethnic background plays a significant role in network formation in educational settings [84].
IV Percolation
The structure of the network determines how robust it is to being disintegrated into small, connected components when links or nodes are removed [85, 86, 87, 88, 89]. In social systems, these removals can stem from communication failures, a limited number of contacts, or other reasons, and the connectivity then determines the spread of ideas, behaviors, or diseases [90, 38].
Removing across-group links—often corresponding to weak ties that bridge otherwise separate communities [91]—can trigger sudden fragmentation, severing pathways across the network. In contrast, losing within-group links, typically associated with strong local ties within groups, tends to undermine cohesion within groups without immediately disrupting global connectivity [92, 93, 94]. This distinction is central to our analysis, where we explicitly differentiate between links that connect groups (the $G_{2}$-layer of our model) and those that bind them internally (the $G_{c}$ layers for $c>2$).
To show how group homophily values $\{h_{c}\}$ influence network behavior, we analyze their effect on connectivity and percolation thresholds. Since our model generates networks with nearly identical degree distributions across different homophily settings, any changes in percolation, such as shifts in the largest connected component or the threshold, reflect structural correlations induced by homophily, not degree variation [95]. This setup isolates the specific role of homophily within and across groups in shaping network robustness and percolation dynamics.
IV.1 Colored Percolation
Figure 3: Percolation Phase Transition in Networks with Different within- and across-homophily Values. Local mixing patterns affect the size of the giant component and the percolation threshold. Panels (a–d) show the percolation transition in the size of the giant component for networks composed of 2- and 6-cliques ($G=G_{2}\cup G_{6}$). Each curve corresponds to a different pair of homophily levels, $h_{2}$ and $h_{6}$, and tracks the growth of the giant component as the red–red link activation probability $\pi_{\mathrm{rr}}$ increases. Link activation settings vary across panels: (a) $\pi_{\mathrm{rb}}=\pi_{\mathrm{bb}}=0.2$; (b) $\pi_{\mathrm{rb}}=0$, $\pi_{\mathrm{bb}}=0.2$; (c) $\pi_{\mathrm{rb}}=0$, $\pi_{\mathrm{bb}}=0$; and (d) $\pi_{\mathrm{rb}}=0.2$, $\pi_{\mathrm{bb}}=0$. All networks are generated with $N=500{,}000$ nodes, average degree six, red node fraction $n_{\mathrm{r}}=0.6$, link composition $\alpha=0.5$, and Coleman index $h=0.3$. Results are averaged over 20 realizations. In (a), all networks remain in the supercritical regime regardless of $\pi_{\mathrm{rr}}$, indicating that an average degree of six is sufficient for global connectivity under moderate homophily. By contrast, in (b) and (d), the percolation order of the curves reverses: the configuration that percolates last in (b) is the first to do so in (d). This reversal highlights how activating within- versus across-group links reshapes the percolation threshold, even when the overall homophily index remains fixed.
In a standard bond percolation process [85], each link in a network $G$ is retained with probability $\phi$ and removed with probability $1-\phi$, independently and uniformly at random. As $\phi$ varies, we track the size of the giant component—the largest connected subgraph—to study how connectivity responds to random removal of links. This process is relevant to various simple dynamics such as epidemics, information spreading, or behavior adoption, where deleted links represent blocked transmission paths. For example, in the case of epidemic spreading, under certain assumptions, the giant component corresponds precisely to the final outbreak size [94, 96]. As the disease becomes more infectious, $\phi$ increases, and the population undergoes a continuous phase transition from a contagion‐free to an endemic state. In more complex scenarios, link removal is not uniform and depends on node attributes. For example, disease transmission probabilities can differ based on age or vaccination status in a partially vaccinated population [97, 98, 99, 100, 101, 102, 103, 104, 14]. In such cases, percolation is no longer governed by a single parameter $\phi$, but by a vector of link-type-specific probabilities. In our colored network model, we remove links based on a vector $(\pi_{\mathrm{rr}},\pi_{\mathrm{rb}},\pi_{\mathrm{bb}})$, where $\pi_{kj}$ is the probability of removing a link between nodes of colors $k$ and $j$ [105]. In the large-network limit, this is equivalent to independently removing each $kj$-type link with probability $\pi_{kj}$. This is similar to the idea of neighbor-induced immunity percolation [106]. By varying these probabilities and observing how the giant component changes, we identify when the network undergoes a phase transition in connectivity.
Now, consider a network where red–blue and blue–blue links remain active with fixed probabilities, while the activation probability of red–red links varies from 0 to 1. In this setup, overall network connectivity is controlled by adjusting the presence of red–red links, $\pi_{\mathrm{rr}}$. Fig. 3(a–d) shows the size of the giant component for networks with the same Coleman index $h=0.3$, but different values of $h_{2}$ and $h_{6}$, which reflect varying levels of across- and within-group homophily.
When red-blue links are partially kept, $\pi_{\mathrm{rb}}>0$, panels (a) and (b), red and blue nodes can form a joint giant component altogether. Even with the same Coleman index $h$, networks with higher across-group homophily $h_{2}$ yield larger giant components at high $\pi_{\mathrm{rr}}$. Therefore, local mixing patterns affect not only the size of the giant component but also when the percolation transition happens, revealing structural effects that measures like the Coleman index miss. When red–blue links are absent, $\pi_{\mathrm{rb}}=0$, percolation can occur only through nodes of the same color, thus the size of the giant component is limited by the number of red or blue nodes. Different combinations of $h_{2}$ and $h_{6}$ modify the percolation threshold. In panel (d), where $\pi_{\mathrm{bb}}=0.2$, the giant component begins growing at the same threshold as in panel (c) ($\pi_{\mathrm{bb}}=0$), but from a higher baseline—indicating that percolation already occurs through blue–blue links, regardless of red-red links. Panel (a) shows that a large cluster emerges regardless of red-red links whenever a fraction of red-blue and blue-blue links are active. In panel (b), percolation thresholds are lower than in panel (c), due to the presence of across-group links. These results show that local mixing patterns affect not only the size of the giant component but also when the percolation transition happens, revealing structural effects that measures like the Coleman index miss.
Figure 4: Within- and across-group Homophily Influences Network Connectivity, Percolation Properties and Epidemic Thresholds in Non-Uniform Ways. Panels (a–c) present heatmaps of the critical percolation value with respect to $h=(h_{2}+h_{4})/2$ and $\Delta h=h_{2}-h_{4}$ for: (a) $\pi_{\mathrm{rr}}=\pi_{\mathrm{bb}}$, $\pi_{\mathrm{rb}}=0.1$; (b) $\pi_{\mathrm{rr}}=\pi_{\mathrm{bb}}$, $\pi_{\mathrm{rb}}=0.5$; (c) $\pi_{\mathrm{rr}}=0.1\pi_{\mathrm{bb}}$ $\pi_{\mathrm{rb}}=0.1,\alpha=0.2$. These panels illustrate how redistributing homophily configurations can either raise or lower the percolation threshold depending on the extent of across-group connectivity $\pi_{rb}$. Panel (d) compares the effect of redistributing homophily across small and large groups in networks with $h=0.5$ and average degree 2. Green regions indicate that emphasizing small-group homophily ($h_{4}=0.1$, $h_{2}=0.9$) leads to a higher percolation threshold than emphasizing large-group homophily ($h_{2}=0.1$, $h_{4}=0.9$). No percolation transition occurs in the top-right white region. Panels (e–g) explore the interaction between homophily and vaccination efficacy with respect to $h=(h_{2}+h_{4})/2$ and $\Delta h=h_{2}h_{4}$. The impact of homophily depends on both within-group efficacy ($f_{v}$) and cross-group efficacy ($f_{I}$): (e) $f_{I}=1$, $f_{v}=0.1$; (f) $f_{I}=0.1$, $f_{v}=0.5$; (g) $f_{I}=0.1$, $f_{v}=1$.
IV.2 Critical percolation value
Analytical solutions for percolation on non-tree-like networks are famously intractable [107, 108]; yet, by isolating within-clique spreading from across-clique behavior, our framework breaks this impasse and provides a closed-form expression for the critical percolation threshold (see Section V.6). Using a multi-type branching process, we show that the percolation vector $\boldsymbol{\pi}=(\pi_{\mathrm{rr}},\pi_{\mathrm{rb}},\pi_{\mathrm{bb}})$ is critical when there exists a non-negative vector $\boldsymbol{v}$ such that
$\boldsymbol{v}\in\ker\big{(}\boldsymbol{B}(\boldsymbol{\pi})-\boldsymbol{I}\big{)},\quad\boldsymbol{v}/|\boldsymbol{v}|\geq 0,$ (9)
where $\boldsymbol{I}$ is the identity matrix and the matrix $\boldsymbol{B}=\boldsymbol{B}(\boldsymbol{\pi})$ (see Eq. (20) in Methods) encodes the expected number of nodes of a given clique type reached from another via percolation. The condition $\det\big{(}\boldsymbol{B}(\boldsymbol{\pi})-\boldsymbol{I}\big{)}=0$ is necessary but not sufficient for criticality. In expression (9), $\ker(\cdot)$ denotes the kernel (null space) of a matrix—i.e., the set of vectors mapped to zero—and $\det(\cdot)$ is the determinant, indicating when the matrix becomes singular.
We explore this condition in networks parametrized by within-group homophily $h_{4}$ and across-group homophily $h_{2}$. Fig. 4(a–b) shows heatmaps of the critical value $\pi_{\mathrm{rr}}=\pi_{\mathrm{bb}}$ for different values of $\pi_{\mathrm{rb}}$. When $\pi_{\mathrm{rb}}$ is low (Fig. 4a), increasing the global homophily expressed by the Coleman index $h$ reduces the critical threshold. When $\pi_{\mathrm{rb}}$ is high (Fig. 4b), increasing $h$ instead raises the threshold. Interestingly, these figures reveal how group homophilies affect connectivity, even when the Coleman index is fixed. We see that level curves for the critical value $\pi_{\mathrm{rr}}=\pi_{\mathrm{bb}}$ are not vertically aligned, meaning that for fixed $h$, distinct combinations of $h_{2}$ and $h_{4}$ yield different percolation thresholds. In particular, when $\pi_{\mathrm{rb}}$ is low, percolation is more easily achieved if $h_{4}>h_{2}$; vice versa, when $\pi_{\mathrm{rb}}$ is high, the percolation threshold is lower if $h_{2}>h_{4}$. The importance of local homophily is more prominent when $\pi_{rr}\neq\pi_{bb}$, as depicted in Fig. 4c. Here, the contour lines are roughly diagonal, indicating that the type of homophily is equally important to the amount of homophily for the percolation threshold. This means that the type of homophily can have a significant impact on the critical value of $\pi_{rr}$, which can range from 0.4 to 0.75 even when the Coleman homophily $h$ remains fixed.
To illustrate the role of homophily within and across groups for the percolation threshold more systematically, we fix $h=0.5$ with $\alpha=0.5$ and compare two configurations: (i) weak across-group homophily ($h_{2}=0.1$) and strong within-group homophily ($h_{4}=0.9$), and (ii) the reverse. We compute the difference in critical values: $\Delta\pi_{\mathrm{rr}}^{*}=\pi_{\mathrm{rr}}^{*}(h_{4}=0.1)-\pi_{\mathrm{rr}}^{*}(h_{2}=0.1)$. Fig. 4d shows that this difference can be either positive or negative, depending on $\pi_{\mathrm{rb}}$ and $\pi_{\mathrm{bb}}$. Increasing the connectivity between groups (higher $\pi_{\mathrm{rb}}$ ) enhances the effectiveness of across-group homophily in maintaining network connectivity. In other words, when many cross-group links are present, networks with stronger across-group homophily achieve percolation at lower thresholds compared to networks with stronger within-group homophily. Thus, stronger homophily within or across groups does not universally improve connectivity. The difference in critical $\pi_{\mathrm{rr}}$ values between networks with the same Coleman index can be as large as 0.8, showing that global homophily alone is insufficient to predict percolation behavior.
IV.3 Epidemic Spread & Vaccination
We demonstrate how multiscale homophily influences epidemic dynamics by applying our percolation framework to contact networks. Homophily affects disease transmission through vaccination behaviors and demographic mixing patterns alike. Age serves as a prime example: empirical contact matrices consistently show strong age-assortative mixing, with children primarily interacting with other children and adults exhibiting distinct interaction patterns depending on age and context [100, 101, 102, 103, 104]. Such structured contacts significantly reshape infection pathways, thereby affecting key epidemiological outcomes including herd immunity thresholds, expected epidemic sizes, and the effectiveness of interventions [14, 25, 26]. Our framework facilitates a detailed analysis of intervention strategies, providing a foundation for improved design and interpretation of public health policies [109]. To illustrate this, we analyze a vaccination scenario involving an imperfect vaccine, categorizing individuals as vaccinated (red) and unvaccinated (blue) nodes, to investigate how homophily shapes disease spread and influences the success of interventions.
With perfect vaccines, only red–red links remain active, and any link involving at least one vaccinated node is removed from the transmission network. However, with imperfect vaccines, vaccinated individuals can still transmit the disease, though at a reduced value, to both vaccinated and unvaccinated individuals compared to the baseline probability $\pi$ between two unvaccinated nodes. If the across-group vaccine efficacy $f_{I}$ reduces transmission between vaccinated and unvaccinated individuals, and the within-group efficacy $f_{v}$ reduces transmission across vaccinated individuals, then the transmission probabilities are $\pi_{\mathrm{rb}}=f_{I}\pi$ and $\pi_{\mathrm{bb}}=f_{v}\pi$, respectively [14].
Fig. 4(e–g) shows that the impact of $f_{I}$ and $f_{v}$ on the critical percolation threshold is non-trivial. As in Fig. 4(a–c), the effect of homophily on the critical value can reverse. In Fig. 4e, increasing either homophily raises the critical threshold, while in Fig. 4g, it lowers it. This mirrors earlier results: for low $\pi_{\mathrm{rb}}$, increasing homophily decreases the critical value. Since $\pi_{\mathrm{rb}}=f_{I}\pi$, the values of $f_{I}$ in these panels explain the observed shift—$\pi_{\mathrm{rb}}=\pi$ in Fig. 4e and $0.1\pi$ in Fig. 4g. Fig. 4f shows that for certain combinations of $f_{I}$ and $f_{v}$, the critical value becomes nearly insensitive to either form of homophily.
In tree-like networks, the required vaccine coverage for herd immunity increases with the Coleman index $h$, and high homophily may render herd immunity unattainable [14]. Our results reveal that layered homophily creates far more complex outcomes. For example, in Fig. 4g, the critical threshold ranges from 0.5 to 0.64 when the Coleman homophily remains fixed at $h=0.5$, a 30% difference, highlighting the influence of homophily structure on epidemic dynamics and intervention outcomes.
V Conclusion & Discussion
We have introduced a flexible framework for capturing homophily across multiple social scales, combining tightly knit subgroups with across-group interactions. By applying a maximum-entropy approach to maximal clique decomposition, our model accurately replicates homophily patterns in diverse real-world datasets, demonstrating how these trends remain hidden when homophily is considered as a single parameter. We further show that homophily may raise or lower percolation thresholds depending on the interaction between within- and across-group percolation, demonstrating that carefully tuned interventions (e.g., targeted vaccination or social distancing) can yield different outcomes. The connectivity insights provided by our framework remain applicable to any spreading phenomenon that can be effectively represented as a percolation process. Thus, our multiscale homophily decomposition provides a principled foundation for analyzing phenomena ranging from information diffusion to countering misinformation by managing link activation within and across groups.
We note that while our model specifically addresses homophilic interactions, it omits other key structural features of social networks, such as degree heterogeneity or degree correlations. This simplification helps isolate the specific effects of group homophily, making the model particularly useful as a null model. Networks that share similar group sizes and homophily levels, yet exhibit additional characteristics like degree heterogeneity, can be compared against our model to determine whether observed differences arise from structural features beyond homophily. Thus, our framework serves to clarify whether empirical patterns are solely driven by homophily or if other structural elements also play a significant role.
Nevertheless, the model can readily incorporate additional constraints on node degrees or hyperedge structures due to its underlying exponential-family formulation, allowing easy customization to empirical data. Furthermore, this flexibility extends naturally to scenarios where homophily parameters evolve over time presents a streamlined yet flexible approach for analyzing multiscale homophily, laying the groundwork for future studies that incorporateltiple attributes per node, capturing interactions between attributes that may result in non-additive effects, such as decreased homophily in one dimension influencing homophily in another [110]. Overall, this work presents a streamlined yet flexible approach for analyzing multiscale homophily, laying the groundwork for future studies that incorporate richer dynamics, degree heterogeneity, and higher-order motifs. This versatility enhances our understanding of how local assortative interactions influence global connectivity across social, biological, and technological networks, prompting further experimental and computational exploration into the broader implications of homophily-like biases in complex systems.
Although our focus has been on social networks, our method can be applied to other systems with “like-likes-like” biases. Polymer or colloidal assemblies, for example, involve monomer or particle species whose binding preferences mirror within-group homophily [111, 112, 113, 111], and neural networks often exhibit connectivity shaped by neuron type or layer [114, 115]. Likewise, in gene regulatory or protein–protein interaction networks, functionally similar nodes tend to cluster [116, 117, 118, 119]. In each of these contexts, our maximum-entropy model could provide a systematic way to assess how local mixing biases influence an entire system.
Our decomposition of homophily reveals that what appears as modest or even absent assortativity at the network level often conceals a spectrum of structured, group-size-dependent mixing patterns. This observation challenges prevailing modeling assumptions and calls for a more elaborate understanding of how homophily operates in social systems and the implications for the various dynamics on them. By capturing this previously hidden structure in homophily patterns, our framework not only offers better alignment with empirical observations but also provides a principled model for investigating the behavior of dynamical processes on social networks.
Acknowledgments
Funding: AKR acknowledges support from the Independent Research Fund Denmark (EliteForsk grant to Sune Lehmann), the Carlsberg Foundation (The Hope Project), and the Villum Foundation (NationScale Social networks). CS is supported by an NWO VENI grant 202.001. MK acknowledges the grant number 349366 from the Research Council of Finland. The simulations presented above were performed using computer resources within the Aalto University School of Science’s “Science-IT” project.
Use of AI tools: We utilized AI language tools to enhance grammar and readability. The authors reviewed and approved all content.
Competing interests: The authors declare no competing interests.
Data and materials availability: All essential simulations and numerical computations are publicly available at [120]. The datasets used in this project were gathered from other research papers. Data collection methods, network details, and attribute distributions are provided in the original papers.